OpenStudy (anonymous):

how do you do the prob 2n^-2??

OpenStudy (anonymous):

What are you asked to do? Convert to positive exponent? Graph?

OpenStudy (anonymous):

convert to positive. sorry.

OpenStudy (lalaly):

\[\frac{2}{n^2}\]because\[x^{-n}=\frac{1}{x^n}\]

OpenStudy (anonymous):

thank you! i totally get it now!

OpenStudy (lalaly):

:)